# Uniform pathwise stability of additive singular SDEs driven by fractional Brownian motion

### v1

@@ 4 Existence and uniqueness of invariant measures
For simplicity, we write $\Psi^{i}_{t}\equiv\Psi(t_{0},t,x,\eta^{i})$. Let $(\tilde{s},\tilde{t})\in\Delta_{[t_{0},T]}$ such that $\tilde{t}-\tilde{s}\leq 1$, and let $(s,t)\in\Delta_{[\tilde{s},\tilde{t}]}$, then
$\Psi_{t}^{1}-\Psi_{t}^{2}-\mathbb{E}^{s}(\Psi_{t}^{1}-\Psi_{t}^{2})$ $=\int_{s}^{t}F(\Psi_{r}^{1})-F(\Psi_{r}^{2})-\mathbb{E}^{s}\left(F(\Psi_{r}^{1})-F(\Psi_{r}^{2})\right)dr$
$\quad+\int_{s}^{t}b(\Psi_{r}^{1})-b(\Psi_{r}^{2})dr-\mathbb{E}^{s}\int_{s}^{t}b(\Psi_{r}^{1})-b(\Psi_{r}^{2})dr.$
Using the Lipschitz property of $F$ and taking the $L_{m}$ norm, we get
$\|\Psi_{t}^{1}-\Psi_{t}^{2}-\mathbb{E}^{s}(\Psi_{t}^{1}-\Psi_{t}^{2})\|_{L_{m}}\leq C\int_{s}^{t}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}dr+C\|\int_{s}^{t}b(\Psi_{r}^{1})-b(\Psi_{r}^{2})dr\|_{L_{m}}.$
Using Proposition 3.4 with $\psi=\Psi-\widetilde{B}^{t_{0}}$, $\phi=\Psi^{2}-\widetilde{B}^{t_{0}}$, $\tau=1/2$, and recalling that $\llbracket\Psi^{1}-\widetilde{B}^{t_{0}}\rrbracket_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[s,t]}L_{1,\infty}}\leq C$ by (4.7), we get that
$\|\Psi_{t}^{1}-\Psi_{t}^{2}-\mathbb{E}^{s}(\Psi_{t}^{1}-\Psi_{t}^{2})\|_{L_{m}}$ $\leq C\sup_{r\in[S,T]}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}(t-s)^{1+H(\gamma-1)}$
$\quad+C\llbracket\Psi^{1}-\Psi^{2}\rrbracket_{\mathcal{C}^{1/2}_{[\tilde{s},\tilde{t}]}L_{m}}(t-s)^{1+H(\gamma-1)}.$
$\llbracket\Psi^{1}-\Psi^{2}\rrbracket_{\mathcal{C}^{1/2}_{[\tilde{s},\tilde{t}]}L_{m}}$ $\leq C\sup_{r\in[S,T]}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}(\tilde{t}-\tilde{s})^{1/2+H(\gamma-1)}$
$\quad+C\llbracket\Psi^{1}-\Psi^{2}\rrbracket_{\mathcal{C}^{1/2}_{[\tilde{s},\tilde{t}]}L_{m}}(\tilde{t}-\tilde{s})^{1/2+H(\gamma-1)}.$
Let $\ell:=(2C)^{\frac{1}{1/2+H(\gamma-1)}}\wedge 1$, then using that $\llbracket\Psi^{1}-\Psi^{2}\rrbracket_{\mathcal{C}^{1/2}_{[\tilde{s},\tilde{t}]}L_{m}}\leq 2[\Psi^{1}-\Psi^{2}]_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[\tilde{s},\tilde{t}]}L_{m}}<+\infty$ (since $\Psi^{1}$ and $\Psi^{2}$ belong to the class $\widetilde{\mathcal{V}}$), we get
$\llbracket\Psi^{1}-\Psi^{2}\rrbracket_{\mathcal{C}^{1/2}_{[\tilde{s},\tilde{t}]}L_{m}}$ $\leq C\sup_{r\in[\tilde{s},\tilde{t}]}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}.$ (4.12)
Next, we compare $\Psi^{1}$ and $\Psi^{2}$ in the supremum norm. Let $t\geq t_{0}$, then
$\Psi_{t}^{1}-\eta^{1}_{t}-(\Psi^{2}_{t}-\eta^{2}_{t})=\int_{t_{0}}^{t}F(\Psi_{r}^{1})-F(\Psi_{r}^{2})dr+\int_{t_{0}}^{t}b(\Psi_{r}^{1})-b(\Psi_{r}^{2})dr.$
$\|\Psi_{t}^{1}-\eta^{1}_{t}-(\Psi^{2}_{t}-\eta^{2}_{t})\|_{L_{m}}\leq\int_{t_{0}}^{t}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}dr+\sum_{k=0}^{N-1}\|\int_{t_{k}}^{t_{k+1}}b(\Psi_{r}^{1})-b(\Psi_{r}^{2})dr\|_{L_{m}}.$
Applying Proposition 3.4 as before and using $\llbracket\Psi^{1}-\widetilde{B}^{t_{0}}\rrbracket_{\mathcal{C}^{1+H(\gamma\wedge 0)}_{[t_{k},t_{k+1}]}L_{1,\infty}}\leq C$, we get that
$\|\Psi_{t}^{1}-\eta^{1}_{t}-(\Psi^{2}_{t}-\eta^{2}_{t})\|_{L_{m}}$ $\leq\int_{t_{0}}^{t}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}dr$
$\quad+C\sum_{k=0}^{N-1}\Big(\sup_{r\in[t_{k},t_{k+1}]}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}(t_{k+1}-t_{k})^{1+H(\gamma-1)}$
$\quad+\llbracket\Psi^{1}-\Psi^{2}\rrbracket_{\mathcal{C}^{1/2}_{[t_{k},t_{k+1}]}L_{m}}(t_{k+1}-t_{k})^{1+H(\gamma-1)}\Big).$
Using (4.12), we have
$\|\Psi_{t}^{1}-\eta^{1}_{t}-(\Psi^{2}_{t}-\eta^{2}_{t})\|_{L_{m}}$ $\leq\int_{t_{0}}^{t}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}dr$
$\quad+C\sum_{k=0}^{N-1}\sup_{r\in[t_{k},t_{k+1}]}\|\Psi_{r}^{1}-\Psi_{r}^{2}\|_{L_{m}}(t_{k+1}-t_{k})^{1+H(\gamma-1)}.$
Hence, we conclude that
$\|\Psi_{t}^{1}-\eta^{1}_{t}-(\Psi^{2}_{t}-\eta^{2}_{t})\|_{L_{m}}$ $\leq C\sum_{k=0}^{N-1}\Big(\sup_{r\in[t_{k},t_{k+1}]}\|\Psi_{r}^{1}-\eta^{1}_{r}-(\Psi^{2}_{r}-\eta^{2}_{r})\|_{L_{m}}$
$\quad+\sup_{r\in[t_{k},t_{k+1}]}\left|\eta^{1}_{r}-\eta^{2}_{r}\right|\Big)(t_{k+1}-t_{k})^{1+H(\gamma-1)}.$
Finally, a Grönwall-type argument shows that
$\|\Psi_{t}^{1}-\eta^{1}_{t}-(\Psi^{2}_{t}-\eta^{2}_{t})\|_{L_{m}}$ $\leq C\sup_{r\in[t_{0},T]}\left|\eta^{1}_{r}-\eta^{2}_{r}\right|,$
which concludes the proof. ∎
Before giving the proof of Theorem 4.3, let us first discuss with the case when $b$ is a smooth function. Here one can solve the deterministic equation $du_{t}=F(u_{t})dt+b(u_{t})dt+df_{t}$ for $t\geq t_{0}$ and for any continuous function $f$, and therefore define the deterministic mapping
$\Gamma:\mathbb{R}_{+}\times\mathbb{R}^{d}\times\mathcal{C}([t_{0},T])\rightarrow\mathbb{R}^{d},$ (4.13)
such that $\Gamma(t_{0},x,f)=u$ (e.g. [12, Proposition 4.3]). In this case, one can check that the process $\omega\mapsto\Gamma(x,\eta+\widetilde{B}^{t_{0}}(\omega))(\omega)$ is a solution to the SDE (4.1), and that the process $\omega\mapsto\Gamma(x,B(\omega))(\omega)$ is a solution to the SDE (1.1). Hence, plugging $\eta=\bar{B}^{t_{0}}$ in the solution of (4.1) gives us the solution to the SDE (1.1).
However, such $\Gamma$ for $b\in\mathcal{C}^{\gamma}$ and $\gamma<0$ does not exist and one needs to rely on smooth approximations of $b$ to get the same result. That is the purpose of Theorem 4.3 which shows how the SDEs (1.1) and (4.1) are related, the intuition still being that plugging $\eta$ to be the history of the fBm in the solution of (4.1) gives the solution of (1.1). The following lemma gives the first part of Theorem 4.3, that is, the joint measurability of $\Psi$ in $\eta$ and $\omega$ up to a modification. This will allow us to consider the map $\omega\mapsto\Psi(t_{0},t,x,\bar{B}^{t_{0}}(\omega))(\omega)$ later in the proof.
Lemma 4.6.
Let $\gamma\in(1-1/(2H),1)$, and let $b$ satisfy Assumption 2.4. Let $F$ satisfy Assumption 2.3 and further assume that $F$ satisfies (2.11) with $\kappa_{2}=0$. For any $x\in\mathbb{R}^{d}$, $T\in\mathbb{R}$, $t_{0}<T$, and any continuous function $\eta:[t_{0},T]\rightarrow\mathbb{R}^{d}$, denote by $(\Psi(t_{0},t,x,\eta))_{t\in[t_{0},T]}$ the unique solution to (4.1) in the class $\widetilde{\mathcal{V}}^{\eta}$ (defined in (4.6)). Then for any $t\in[t_{0},T]$, $(\eta,\omega)\mapsto\Psi(t_{0},t,x,\eta)(\omega)$ has a jointly measurable modification, which we still call $\Psi$.
Proof.
Let $T\in\mathbb{R}$, $t_{0}<T$ and $x\in\mathbb{R}^{d}$. We take $(b^{n})_{n\in\mathbb{N}}$ a sequence of smooth functions converging to $b$ in $\mathcal{C}^{\gamma}$ such that $\sup_{n\in\mathbb{N}}\|b^{n}\|_{\mathcal{C}^{\gamma}}\leq\Xi$, we write $\Psi^{n}$ for the solution of (4.1) with drift $b^{n}$.
Let $t\in[t_{0},T]$. For any continuous function $\eta:[t_{0},T]\rightarrow\mathbb{R}^{d}$, define $f(\eta,\omega):=\Psi(t_{0},t,x,\eta)(\omega)$. Then for each $\eta$, $f(\eta,\cdot)$ is a measurable function, since $\Psi$ is measurable in $\omega$ as the limit of $\Psi^{n}$ by (4.8). Moreover, for any sequence $\eta_{n}$ that converges to $\eta$ in the space of continuous functions over $t_{0},T]$ with respect to the supremum norm, we have that $\Psi(t_{0},t,x,\eta_{n})$ converges to $\Psi(t_{0},t,x,\eta)$ in probability by Proposition 4.5. So it follows that $f$ is stochastically continuous and one can apply [36, Lemma 4.31] to conclude that $f$ has a jointly measurable modification. ∎
We are now ready to give the proof of Theorem 4.3.
Proof of Theorem 4.3..
Let $T\in\mathbb{R}$, $t_{0}<T$ and $x\in\mathbb{R}^{d}$. We take $(b^{n})_{n\in\mathbb{N}}$ a sequence of smooth functions converging to $b$ in $\mathcal{C}^{\gamma}$ such that $\sup_{n\in\mathbb{N}}\|b^{n}\|_{\mathcal{C}^{\gamma}}\leq\Xi$, we write $\Psi^{n}$ for the solution of (4.1) with drift $b^{n}$, and we write $\Gamma^{n}$ for the deterministic function defined in (4.13) with drift $b^{n}$. By Proposition 3.10, we know that up to some subsequence, for all $t\in[t_{0},T]$, $\Gamma^{n}(t_{0},x,B)(t)$ converges to $X^{x}_{t}$ almost surely. Since the subsequence depends only on $T$, we assume without any loss of generality that for all $t$, $\Gamma^{n}(t_{0},x,B)(t)$ converges almost surely to $X^{x}_{t}$.
Let $t\in[t_{0},T]$. For any $n\in\mathbb{N}$ and any continuous function $\eta:[t_{0},T]\rightarrow\mathbb{R}^{d}$, define $f^{n}(\eta,\omega):=|\Psi^{n}(t_{0},t,x,\eta)(\omega)-\Psi(t_{0},t,x,\eta)(\omega)|\wedge 1$. Then, Lemma 4.6 gives the joint measurability of $\Psi$ (up to a modification that we still call $\Psi$). Moreover, since $\Psi^{n}$ is jointly measurable, we also get that $f^{n}$ has a jointly measurable modification which we still denote by $f^{n}$.
Moreover, for any $\eta$, $f^{n}(\eta,\cdot)$ is independent of $\bar{B}^{t_{0}}$ (the solution of (4.1) is independent of the history of the fBm since $\widetilde{B}^{t_{0}}$ is independent of $\bar{B}^{t_{0}}$). Then by Lemma A.3 (where one takes $\mathcal{G}$ to be the sigma algebra generated by $\widetilde{B}^{t_{0}}$), we get that
$\mathbb{E}f^{n}(\bar{B}^{t_{0}}(\cdot),\cdot)=\mathbb{E}g^{n}(\bar{B}^{t_{0}}),$
where $g^{n}(\eta)=\mathbb{E}f^{n}(\eta,\cdot)$. Moreover, by (4.8), we have $\sup_{\eta}g^{n}(\eta)=\sup_{\eta}\mathbb{E}|\Psi^{n}(t_{0},t,x,\eta)-\Psi(t_{0},t,x,\eta)|\leq C\|b^{n}-b\|_{\mathcal{C}^{\gamma-1}}$. Hence, it follows that
$\mathbb{E}f^{n}(\bar{B}^{t_{0}}(\cdot),\cdot)=\mathbb{E}|\Psi^{n}(t_{0},t,x,\bar{B}^{t_{0}})-\Psi(t_{0},t,x,\bar{B}^{t_{0}})|\leq C\|b^{n}-b\|_{\mathcal{C}^{\gamma-1}}.$
Now notice that $\Psi^{n}(t_{0},t,x,\bar{B}^{t_{0}})=\Gamma^{n}(t_{0},x,B)(t)$, hence by letting $n\rightarrow\infty$ and recalling that $\Gamma^{n}(t_{0},x,B)(t)$ converges almost surely to $X^{x}_{t}$, we conclude that $\mathbb{E}|X^{x}_{t}-\Psi(t_{0},t,x,\bar{B}^{t_{0}})|=0$. ∎
Existence of an invariant measure.
In [19], the author shows that if the evolution $\xi$, started from the generalised initial condition $\delta_{x}\times\mathbf{W}$ admits a Lyapunov function, then there exists an invariant measure $\mu$. This is achieved by using a Krylov-Bogoliubov argument, proving tightness of the law of the evolution, and then using the Feller property to pass to the limit. Let us therefore prove the Feller property and show that the evolution $\xi$ admits a Lyapunov function.
Recall the definition of $\xi$ and the semigroup associated to $\xi$ defined in (4.4) and (4.5). The Feller property of the evolution $\xi$ comes down to the continuity of $\mathbb{E}f(\xi(t_{0},t,x,w))$ for bounded continuous $f$ with respect to $x$ and $w$. This follows from first the continuity of the operator $\mathbf{A}^{t_{0}}$ and then from both (4.9) and Proposition 4.5 which state continuity in moments of $\Psi^{t_{0}}$ with respect to $x$ and $\mathbf{A}^{t_{0}}(w)$.
Let $x\in\mathbb{R}^{d}$. Define $\bar{W}=(W_{t+t_{0}})_{t\leq 0}$. Then $\bar{W}$ has law $\mathbf{W}$ and $\mathbf{A}^{t_{0}}(\bar{W})=\bar{B}^{t_{0}}$. Hence, we have $\mathbb{E}|\Psi^{t_{0}}(t_{0},t,x,\mathbf{A}^{t_{0}}(\bar{W}))|^{2}=\mathbb{E}|\Psi^{t_{0}}(t_{0},t,x,\bar{B}^{t_{0}})|^{2}$. Using (3.1) and (4.10), we get that
$\mathbb{E}|\Psi^{t_{0}}(t_{0},t,x,\mathbf{A}^{t_{0}}(\bar{W}))|^{2}=\mathbb{E}|X_{t}|^{2}\leq C.$
It follows that $x\mapsto|x|^{2}$ is a Lyapunov function for the evolution $\xi$ started from the generalised initial condition $\delta_{x}\times\mathbf{W}$. Therefore, we conclude the existence of an invariant measure.
Uniqueness of the invariant measure via the exponential decay.
We prove now the uniqueness of the invariant measure. Let $m\in[2,\infty)$, we actually prove a stronger result, which is an exponential contraction of solutions under the assumption that the Besov-Hölder norm of $\|b\|_{\mathcal{C}^{\gamma}}$ is small enough so that
$-\kappa_{1}+\mathbf{M}(m,\gamma,H,L_{F},\Xi)\|b\|_{\mathcal{C}^{\gamma}}<0,$
where recall that $\kappa_{1}$ is the dissipation constant in (2.11) and $\mathbf{M}(m,\gamma,H,L_{F})$ is the constant from (4.9). Let $\mu,\nu$ denote two invariant measures associated to (4.1). Let $(Z_{1},\bar{W}_{1})$, $(Z_{2},\bar{W}_{2})$ be two random variables defined on some probability space $(\Omega^{\prime},\mathcal{F}^{\prime},\mathbb{P}^{\prime})$ such that their respective laws are $\mu$ and $\nu$. Since the marginals of $\mu$ and $\nu$ over $\mathcal{H}_{H}$ are the same, using the gluing lemma ([39]), one can construct a triple $(X_{0},W,Y_{0})$ such that $(X_{0},W)$ has law $\mu$ and $(Y_{0},W)$ has law $\nu$.
On $\Omega^{\prime}\times\Omega$, we consider the evolutions
$\left(\xi(t_{0},t,X_{0}(\omega^{\prime}),W(\omega^{\prime}))(\omega)\right)_{t\geq t_{0}}\ \text{and}\ \left(\xi(t_{0},t,Y_{0}(\omega^{\prime}),W(\omega^{\prime}))(\omega)\right)_{t\geq t_{0}},$
@@
which by definition of the invariant measure, have respective laws $\mu$ and $\nu$. Therefore, equipping the space $\mathbb{R}^{d}\times\mathcal{H}_{H}$ with the norm
@@
$\mathcal{W}_{m}(\mu,\nu)$ $\leq\big(\mathbb{E}\|\xi(t_{0},t,X_{0},W)-\xi(t_{0},t,Y_{0},W)\|^{m}\big)^{1/m}$
$=\big(\mathbb{E}_{\mathbb{P}^{\prime}}f(X_{0},Y_{0},W)\big)^{1/m},$
where
$f(x,y,w):=\mathbb{E}_{\mathbb{P}}|\Psi(t_{0},t,x,\mathbf{A}^{t_{0}}(w))-\Psi(t_{0},t,y,\mathbf{A}^{t_{0}}(w))|^{m}.$
Using (4.9), it follows that
$\mathcal{W}_{m}(\mu,\nu)$ $\leq\exp\{-\kappa_{1}(t-t_{0})+\mathbf{M}(m,\gamma,H,L_{F},\Xi)\|b\|_{\mathcal{C}^{\gamma}}(t-t_{0})\}(\mathbb{E}_{\mathbb{P^{\prime}}}|X_{0}-Y_{0}|^{m})^{1/m}.$
Letting $t$ go to $\infty$, we conclude that $\mathcal{W}_{m}(\mu,\nu)=0$, so $\mu=\nu$.
@@ Appendix A Auxiliary results
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $E$ be a topological space space endowed with its Borel $\sigma$–algebra $\mathcal{B}(E)$. Let $X:\Omega\to E$ be an $E$-valued random variable, and $\mathcal{F}_{X}:=\sigma(X)$. Let $\mathcal{G}\subseteq\mathcal{F}$ be a sigma algebra independent of $\sigma(X)$. Let $f:E\times\Omega\to\mathbb{R}$ be a bounded $(\mathcal{B}(E)\otimes\mathcal{G})$–measurable function, then
$\mathbb{E}f\big(X(\cdot),\cdot\big)=\mathbb{E}g_{f}\big(X(\cdot)\big),$ (A.4)
@@
Let $A\in\mathcal{B}(E)$ and $B\in\mathcal{G}$ and set $U:=A\times B$. Assume first that $f(x,\omega):=\mathds{1}_{A}(x)\,\mathds{1}_{B}(\omega)$. By the tower property of conditional expectation and the independence between $\mathds{1}_{B}$ and $X$, we obtain
$\mathbb{E}f(X(\cdot),\cdot)=\mathbb{E}\mathds{1}_{A}\bigl(X\bigr)\,\mathds{1}_{B}=\mathbb{E}\,\mathbb{E}\bigl[\mathds{1}_{A}\bigl(X\bigr)\,\mathds{1}_{B}\mid\mathcal{F}_{X}\bigr]=\mathbb{E}\bigl[\mathds{1}_{A}(X)\,\mathbb{E}[\mathds{1}_{B}]\bigr]=\mathbb{E}g_{f}(X(\cdot),\cdot).$
@@
$\mathcal{H}:=\Bigl\{k:\,\text{bounded, }(\mathcal{B}(E)\otimes\mathcal{G})\text{-measurable and }\mathbb{E}k(X(\cdot),\cdot)=\mathbb{E}g_{k}(X(\cdot),\cdot)\}$
@@
Hence, by the functional monotone-class theorem (e.g. [13, Theorem 5.2.2]), $\mathcal{H}$ contains all bounded functions that are $\sigma(\mathscr{A})$-measurable. Since $\sigma(\mathscr{A})=\mathcal{B}(E)\otimes\mathcal{G}$, we conclude that every bounded $\mathcal{B}(E)\times\mathcal{G}$-measurable $k$ lies in $\mathcal{H}$. This completes the proof. ∎

### v2

@@ 4 Markovian structure and invariant measures
we consider for any $t\geq t_{0}$ the random variables
$\Theta(t_{0},t,X_{0},\bar{W})\ \text{and}\ \Theta(t_{0},t,Y_{0},\bar{W}).$
By definition of the invariant measure, for any bounded measurable function $f:\mathbb{R}^{d}\times\mathcal{H}_{H}\rightarrow\mathbb{R}$, we have for any $t\geq t_{0}$
$\widetilde{\mathbb{E}}f(X_{0},\bar{W})$ $=\mathbb{E}_{\mathbb{P}^{\prime}}\left(\mathbb{E}_{\mathbb{P}}f(\Theta(t_{0},t,x,w))\right)_{(x,w)=(X_{0},\bar{W})}$
$\widetilde{\mathbb{E}}f(Y_{0},\bar{W})$ $=\mathbb{E}_{\mathbb{P}^{\prime}}\left(\mathbb{E}_{\mathbb{P}}f(\Theta(t_{0},t,y,w))\right)_{(y,w)=(Y_{0},\bar{W})}.$
@@
Hence, using Lemma A.3, we deduce that $\Theta(t_{0},t,X_{0},\bar{W})$ and $\Theta(t_{0},t,Y_{0},\bar{W})$ have respective laws $\mu$ and $\nu$ for any $t$. Therefore, equipping the space $\mathbb{R}^{d}\times\mathcal{H}_{H}$ with the norm
@@
$\mathcal{W}_{m}(\mu,\nu)$ $\leq\big(\widetilde{\mathbb{E}}\|\Theta(t_{0},t,X_{0},\bar{W})-\Theta(t_{0},t,Y_{0},\bar{W})\|^{m}\big)^{1/m}$
$=\big(\widetilde{\mathbb{E}}|X(t_{0},t,x,\mathbf{A}^{t_{0}}(\bar{W}))-X(t_{0},t,y,\mathbf{A}^{t_{0}}(\bar{W}))|^{m}\big)^{1/m}.$
Using Lemma 4.8, (4.12) and (4.22), it follows that
$\mathcal{W}_{m}(\mu,\nu)$ $\leq C\exp\{-\kappa_{1}(t-t_{0})+\mathbf{M}(m,\gamma,H,L_{F},\Xi)\|b\|_{\mathcal{C}^{\gamma}}(t-t_{0})\}(\mathbb{E}_{\mathbb{P^{\prime}}}|X_{0}-Y_{0}|^{m})^{1/m}.$
Taking $m=2$ and letting $t$ go to $\infty$, we conclude that $\mathcal{W}_{2}(\mu,\nu)=0$, so $\mu=\nu$. Choosing $(X_{0},\bar{W})$ to have law $\mu_{0}$ rather than $\mu$, one can run the same argument to show that
$\mathcal{W}_{m}(\mathcal{L}(\Theta(t_{0},t,X_{0},\mathbf{A}^{t_{0}}(\bar{W})),\mu)\leq C\exp\{\left(-\kappa_{1}+\mathbf{M}(m,\gamma,H,L_{F},\Xi)\|b\|_{\mathcal{C}^{\gamma}}\right)(t-t_{0})\},$
and use Lemma 4.8 to get (4.13).
@@ Appendix A Auxiliary results
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and let $E$ be a topological space space endowed with its Borel $\sigma$-algebra $\mathcal{B}(E)$. Let $J:\Omega\to E$ be an $E$-valued random variable, and $\mathcal{F}_{J}:=\sigma(J)$. Let $\mathcal{G}\subseteq\mathcal{F}$ be a sigma algebra independent of $\sigma(J)$. Let $f:E\times\Omega\to\mathbb{R}$ be a bounded $(\mathcal{B}(E)\otimes\mathcal{G})$–measurable function, then
$\mathbb{E}f\big(J(\cdot),\cdot\big)=\mathbb{E}g_{f}\big(J(\cdot)\big),$ (A.4)
@@
Let $A\in\mathcal{B}(E)$ and $B\in\mathcal{G}$ and set $U:=A\times B$. Assume first that $f(x,\omega):=\mathds{1}_{A}(x)\,\mathds{1}_{B}(\omega)$. By the tower property of conditional expectation and the independence between $\mathds{1}_{B}$ and $J$, we obtain
$\mathbb{E}f(Y(\cdot),\cdot)=\mathbb{E}\mathds{1}_{A}\bigl(J\bigr)\,\mathds{1}_{B}=\mathbb{E}\,\mathbb{E}\bigl[\mathds{1}_{A}\bigl(J\bigr)\,\mathds{1}_{B}\mid\mathcal{F}_{J}\bigr]=\mathbb{E}\bigl[\mathds{1}_{A}(J)\,\mathbb{E}[\mathds{1}_{B}]\bigr]=\mathbb{E}g_{f}(J(\cdot),\cdot).$
@@
$\mathcal{H}:=\Bigl\{k:\,\text{bounded, }(\mathcal{B}(E)\otimes\mathcal{G})\text{-measurable and }\mathbb{E}k(J(\cdot),\cdot)=\mathbb{E}g_{k}(J(\cdot),\cdot)\}$
@@
Hence, by the functional monotone-class theorem (e.g. [15, Theorem 5.2.2]), $\mathcal{H}$ contains all bounded functions that are $\sigma(\mathscr{A})$-measurable. Since $\sigma(\mathscr{A})=\mathcal{B}(E)\otimes\mathcal{G}$, we conclude that every bounded $\mathcal{B}(E)\times\mathcal{G}$-measurable $k$ lies in $\mathcal{H}$. This completes the proof. ∎